OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12 answer choices 3y + 12 -3y + 12 3y - 12 -3y - 12

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

"isolate x in the second equation" means move x on one side of = n everything else on the other...

OpenStudy (anonymous):

i know but then it asks about the 1st equation

OpenStudy (superdavesuper):

dats AFTER u isolate x; then u put it in the 1st eqn. so the Q itself is asking abt the 2nd eqn only.

OpenStudy (anonymous):

ok so im guessing its C

OpenStudy (superdavesuper):

nope - rearrange 2nd eqn until u get x=....

OpenStudy (anonymous):

B is my 2nd guess cuz of the positive 12

OpenStudy (superdavesuper):

bingo :)

OpenStudy (anonymous):

yay, thanks ! can you help me with another one ?

OpenStudy (superdavesuper):

only if its not 2 complicated :P

OpenStudy (anonymous):

system A 3x + 2y = 3 –2x – 8y = –1 system b –x - 14y = 1 –2x - 8y = –1 Which of the following statements is correct about the two systems of equations? The value of x for System B will be one–third of the value of x for System A because the coefficient of x in the first equation of System B is one third times the coefficient of x in the first equation of System A. The value of x for System A will be equal to the value of y for System B because the first equation of System B is obtained by adding –4 to the first equation of System A and the second equations are identical. They will have the same solution because they represent the same lines when plotted on the coordinate axes. They will have the same solution because the first equation of System B is obtained by adding the first equation of System A to 2 times the second equation of System A.

OpenStudy (superdavesuper):

AHHHHH i thought i said not 2 complicated...

OpenStudy (anonymous):

you hot girl tee hee

OpenStudy (anonymous):

ok, ill get another question, A system of equations is shown below: 5x + 2y = 3 (equation 1) 2x – 3y = 1 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof? Show that the solution to the system of equations 7x – y = 4 and 2x – 3y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 2x + 5y = 3 and 3x – 2y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 9x + 4y = 5 and 7x – y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations –4x + 9y = 5 and 2x – 3y = 1 is the same as the solution to the given system of equations

OpenStudy (anonymous):

your hot tee hee

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

Done? :)

OpenStudy (anonymous):

not quite, i need help with one more.

OpenStudy (superdavesuper):

hahahaa but some1 else is helping w dat already :)

OpenStudy (anonymous):

@superdavesuper Midge baked 6 cookies and 4 brownies. Midge only has enough ingredients to bake at most 25 cookies or brownies total. Let x represent the number of more cookies and y represent the number of more brownies that Midge can bake. Which of the following graphs best represents the relationship between x and y?

OpenStudy (superdavesuper):

OK a few things from the Q: at most 25 cookies or brownies...but already baked 6 cookies AND 4 brownies. So how much ingredients are left? Take a guess plz? :)

OpenStudy (anonymous):

15

OpenStudy (superdavesuper):

Good :) then which graph u think is related to 15?

OpenStudy (anonymous):

the first one.

OpenStudy (superdavesuper):

yup yup u are getting GOOD at math @yeah_i_suck_at_math lol

OpenStudy (anonymous):

yay, thanks